isaac151:

An investment of 40,000 was made by a business club. The investment was split into three parts and lasted for one year. The first part of the investment earned 8% interest, the second 6%, and the third 9%. Total interest from the investments was 3060. The interest from the first investment was 2 times the interest from the second. Find the amounts of the three parts of the investment.

supie:

\[3060 = .08(2x) + .06(x) + .09(40000 - 3x)\]\[06000 = 8(2x) + 6(x) + 9(40000 - 3x)\]\[306000 = 16x + 6x + 360000 - 27x\]\[306000 = -5x + 360000\]\[5x = 54000\]\[x = 10500\]

supie:

8% = $21000 6% = $10500 9% = $8500